# Uncertainty-aware No-Reference Point Cloud Quality Assessment

### v1

@@ Abstract
The evolution of compression and enhancement algorithms necessitates an accurate quality assessment for point clouds. Previous works consistently regard point cloud quality assessment (PCQA) as a MOS regression problem and devise a deterministic mapping, ignoring the stochasticity in generating MOS from subjective tests. Besides, the viewpoint switching of 3D point clouds in subjective tests reinforces the judging stochasticity of different subjects compared with traditional images. This work presents the first probabilistic architecture for no-reference PCQA, motivated by the labeling process of existing datasets. The proposed method can model the quality judging stochasticity of subjects through a tailored conditional variational autoencoder (CVAE) and produces multiple intermediate quality ratings. These intermediate ratings simulate the judgments from different subjects and are then integrated into an accurate quality prediction, mimicking the generation process of a ground truth MOS. Specifically, our method incorporates a Prior Module, a Posterior Module, and a Quality Rating Generator, where the former two modules are introduced to model the judging stochasticity in subjective tests, while the latter is developed to generate diverse quality ratings. Extensive experiments indicate that our approach outperforms previous cutting-edge methods by a large margin and exhibits gratifying cross-dataset robustness.
@@ 1 Introduction
Point clouds are essential representations of 3D scenes and have found wide applications in various fields, such as autonomous driving and computer vision. Due to the unaffordable complexity and size of raw point clouds, compression techniques are commonly used to reduce relevant processing and storage costs. However, compression can lead to visual quality degradation. Therefore, accurate measurement of point cloud quality is crucial to assess the fidelity of processed data and serves as a foundation for benchmarking and improving point cloud processing algorithms. Similar to the relevant definitions for image quality assessment, point cloud quality assessment (PCQA) can be categorized into three factions: full-reference PCQA (FR-PCQA), reduced-reference PCQA (RR-PCQA), and no-reference PCQA (NR-PCQA), depending on the availability of reference point clouds. As pristine point clouds are often transparent in typical client applications, NR-PCQA has received significant attention and exploration efforts.
@@
Figure 2: Overall architecture of the proposed method. Our method comprises a Prior Module ($\text{PM}_{\text{prior}}$), a Posterior Module ($\text{PM}_{\text{post}}$), and a Quality Rating Generator (QRG). The $\text{PM}_{\text{prior}}$ and $\text{PM}_{\text{post}}$ encode the $N_{v}$ projections from point cloud $I$, or along with its MOS $M$ for $\text{PM}_{\text{post}}$, into a Gaussian variable $z$ to model the stochasticity in subjective tests, while the QRG infers the quality of point cloud $I$ by utilizing the sampled stochastic feature and projections. Note that the two QRGs share network parameters. Solid lines indicate functionality in both the training and testing phases, while dashed lines indicate only working in the training stage.
@@ 2.1 Point Cloud Quality Assessment
The surging popularity of point clouds has sparked significant interest in PCQA. The development of PCQA has drawn inspiration from the advancements in image quality assessment, leading to categorizing methods into FR-PCQA, RR-PCQA, and NR-PCQA concerning their reliance on pristine point clouds. Early interest in FR-PCQA emerges from MPEG, with a primary focus on evaluating and optimizing point cloud compression algorithms. For example, point-to-point error Mekuria et al. (2016) and point-to-plane error Tian et al. (2017) are proposed to measure the geometry quality of point clouds and lay a solid foundation for subsequent investigations Alexiou and Ebrahimi (2018); Javaheri et al. (2020a); Meynet et al. (2019); Yang et al. (2020a). Inspired by the widely-used metrics in images (i.e., PSNR and SSIM), Torlig et alTorlig et al. (2018) and Alexiou et alAlexiou and Ebrahimi (2020) design the point cloud-based versions, respectively, namely PSNR-yuv and PointSSIM. Yang et alYang et al. (2020b) infer the perceptual quality of point clouds using graph signal gradient and construct local graph representations. PCQM Meynet et al. (2020) considers both geometry and color distortions and utilizes a weighted combination to calculate a comprehensive quality description of point clouds. Recently, a novel distortion quantification method is devised by Yang et alYang et al. (2021) to model point cloud quality via multi-scale potential energy discrepancy. Apart from the aforementioned FR-PCQA works, some notable RR-PCQA methods Liu et al. (2021a); Zhou et al. (2023); Liu et al. (2022b) are also proposed and enrich the research on PCQA.
@@ 2.2 Generative Model
In the quality assessment field, generative adversarial networks (GAN) as another type of generative model have been exploited by some works Lin and Wang (2018); Ma et al. (2020); Zhu et al. (2021); Ren et al. (2018); Yang et al. (2020c) to assess image quality. Moreover, most of these works employ the framework of GANs to generate pseudo-reference images or for transfer learning purposes. To the best of our knowledge, no work explores the utilization of VAEs or CVAEs in the context of quality assessment.
@@ 4.2 Datasets & Evaluation Metrics
Evaluation Metrics. We leverage four widely adopted evaluation metrics to quantify the level of agreement between predicted quality scores and ground truth MOSs. These metrics encompass the Spearman Rank Correlation Coefficient (SRCC), Pearson Linear Correlation Coefficient (PLCC), Kendall’s Rank Correlation Coefficient (KRCC), and Root Mean Squared Error (RMSE). Since there can be value range misalignment between predicted results and ground truth MOSs, we follow the previous method Zhang et al. (2023) and introduce a common four-parameter logistic function Antkowiak et al. (2000) to align their range. By employing the aforementioned four metrics, we can obtain convincing benchmarking results.
@@ 4.5 Ablation Studies
To study the effectiveness of our design details, we tune our model on the WPC2.0 dataset and analyze the performance changes.

### v2

# Stochasticity-aware No-Reference Point Cloud Quality Assessment

@@ Abstract
The evolution of point cloud processing algorithms necessitates an accurate assessment for their quality. Previous works consistently regard point cloud quality assessment (PCQA) as a MOS regression problem and devise a deterministic mapping, ignoring the stochasticity in generating MOS from subjective tests. This work presents the first probabilistic architecture for no-reference PCQA, motivated by the labeling process of existing datasets. The proposed method can model the quality judging stochasticity of subjects through a tailored conditional variational autoencoder (CVAE) and produces multiple intermediate quality ratings. These intermediate ratings simulate the judgments from different subjects and are then integrated into an accurate quality prediction, mimicking the generation process of a ground truth MOS. Specifically, our method incorporates a Prior Module, a Posterior Module, and a Quality Rating Generator, where the former two modules are introduced to model the judging stochasticity in subjective tests, while the latter is developed to generate diverse quality ratings. Extensive experiments indicate that our approach outperforms previous cutting-edge methods by a large margin and exhibits gratifying cross-dataset robustness. Codes are available at https://git.openi.org.cn/OpenPointCloud/nrpcqa.
@@ 1 Introduction
Point clouds are essential representations of 3D scenes and have found wide applications in various fields Qi et al. (2017); Fan et al. (2024). Due to the unaffordable complexity and size of raw point clouds, compression techniques are commonly used to reduce relevant processing and storage costs. However, compression can lead to visual quality degradation. Therefore, accurate measurement of point cloud quality is crucial to assess the fidelity of processed data and serves as a foundation for benchmarking and improving point cloud processing algorithms. Similar to the definitions for image quality assessment, point cloud quality assessment (PCQA) can be categorized into three factions: full-reference PCQA (FR-PCQA), reduced-reference PCQA (RR-PCQA), and no-reference PCQA (NR-PCQA), depending on the availability of reference data. As pristine data are often transparent in typical client applications, NR-PCQA has received significant attention and exploration efforts.
@@
Figure 2: Overall architecture of the proposed method. Our method comprises a Prior Module ($\text{PM}_{\text{prior}}$), a Posterior Module ($\text{PM}_{\text{post}}$), and a Quality Rating Generator (QRG). Note that the two QRGs share network parameters. Solid lines indicate functionality in both the training and testing phases, while dashed lines indicate only working in the training stage.
@@ 2.1 Point Cloud Quality Assessment
The surging popularity of point clouds has sparked significant interest in PCQA. The development of PCQA has drawn inspiration from the advancements in image quality assessment, leading to categorizing methods into FR-PCQA, RR-PCQA, and NR-PCQA concerning their reliance on pristine point clouds. Early interest in FR-PCQA emerges from MPEG, with a primary focus on evaluating and optimizing point cloud compression algorithms. For example, point-to-point error Mekuria et al. (2016) and point-to-plane error Tian et al. (2017) are proposed to measure the geometry quality of point clouds and lay a solid foundation for subsequent investigations Alexiou and Ebrahimi (2018); Javaheri et al. (2020a); Meynet et al. (2019); Yang et al. (2020a). Inspired by the widely-used metrics in images (i.e., PSNR and SSIM), Torlig et alTorlig et al. (2018) and Alexiou et alAlexiou and Ebrahimi (2020) design the point cloud-based versions, respectively, namely PSNR-yuv and PointSSIM. Yang et alYang et al. (2020b) infer the perceptual quality of point clouds using graph signal gradient and construct local graph representations. PCQM Meynet et al. (2020) considers both geometry and color distortions and utilizes a weighted combination to calculate a comprehensive quality description of point clouds. Recently, a novel distortion quantification method is devised by Yang et alYang et al. (2021) to model point cloud quality via multi-scale potential energy discrepancy. Apart from the aforementioned FR-PCQA works, some notable RR-PCQA methods Liu et al. (2021a, 2022b) are also proposed and enrich the research on PCQA.
@@ 2.2 Generative Model
In the quality assessment field, generative adversarial networks (GAN) as another type of generative models have been exploited by some works Lin and Wang (2018); Ma et al. (2020); Zhu et al. (2021); Ren et al. (2018); Yang et al. (2020c) to assess image quality. Moreover, most of these works employ the framework of GANs to generate pseudo-reference images or for transfer learning purposes. To the best of our knowledge, no work explores the utilization of VAEs or CVAEs in the context of quality assessment.
@@ 4.2 Datasets & Evaluation Metrics
Evaluation Metrics. We leverage four widely adopted evaluation metrics to quantify the level of agreement between predicted quality scores and ground truth MOSs. These metrics encompass the Spearman Rank Correlation Coefficient (SRCC), Pearson Linear Correlation Coefficient (PLCC), Kendall’s Rank Correlation Coefficient (KRCC), and Root Mean Squared Error (RMSE). Since there can be value range misalignment between predicted results and ground truth MOSs, we follow the previous method Zhang et al. (2023) and introduce a common four-parameter logistic function to align their range. By employing the aforementioned four metrics, we can obtain convincing benchmarking results.
@@ 4.5 Ablation Studies
To study the effectiveness of our designs , we tune our model on the WPC2.0 dataset and analyze the performance changes.
@@ 5 Conclusion
Acknowledgments
This work was supported by The Major Key Project of PCL (PCL2024A02), Natural Science Foundation of China (62271013, 62031013), Guangdong Provincial Key Laboratory of Ultra High Definition Immersive Media Technology (2024B1212010006), Guangdong Province Pearl River Talent Program (2021QN020708), Shenzhen Science and Technology Program (JCYJ20240813160202004, JCYJ20230807120808017), Guangdong Basic and Applied Basic Research Foundation (2024A1515010155).